# The magnitude and spectral geometry

### v2

@@ 1.1. Acknowledgements
We thank Daniel Grieser, Gerd Grubb and Mark Meckes for useful comments on this article and Tony Carbery, Daniel Grieser, Gerd Grubb, Tom Leinster, Rafe Mazzeo, Mark Meckes, Niels Martin Møller, Grigori Rozenblum, Jan-Philip Solovej and Simon Willerton for fruitful discussions.
@@ 2. Overview and main results
    Analogous to Kac’ famous question [19], one can “magnitude the shape of a drum” for balls: if $X$ is a smooth, compact domain in $\mathbb{R}^{n}$, $B$ is a ball and $\mathcal{M}_{X}(R)=\mathcal{M}_{B}(R)+o(R^{n-1})$, then $X$ is isometric to $B$. See Subsection 3.3.
@@
Suppose that $M\subseteq\mathbb{R}^{N}$ is a compact $n-1$-dimensional submanifold, $T>0$ a parameter and $X_{T}=M\times[0,T]\subset\mathbb{R}^{N+1}$ is the cylinder on $M$ of height $T$. The space $X_{T}$ is a compact, $n$-dimensional manifold with boundary, depicted in Figure 1(a). Equipping $X_{T}$ with the subspace distance, Proposition 3.9 relates the magnitude of $X_{T}$ with the magnitude of the base $M$:
@@
Indeed, the volume of $X_{r}$ is $\pi(\sinh(2r)-2r)$, the area of $\partial X_{r}$ is $2\pi(\cosh(2r)-1)$, $s=-6$ in real hyperbolic $3$-space and $H=H_{0}(r)$ is the mean curvature of the boundary of the ball, which only depends on $r$ by rotational invariance.
@@
where $H_{\mathrm{d}^{2}}(x,\cdot)$ is a Riemannian metric on $X$. In local coordinates $C^{j}$ is a symmetric $j$-form in $x-y$. The Taylor expansion (3) translates into an expansion of the symbol of the pseudodifferential operator $Q_{X}$, see Theorem 5.3, where terms for larger $j$ contribute to terms of the symbol of lower order jointly in $\xi$ and $R$.
@@
The analysis in Section 5 for a closed manifold illustrates the general approach taken in this article. For a manifold with boundary the inverse $Q_{X}(R)^{-1}$ decomposes into the previously studied interior part and a new boundary contribution, as described in Theorem 6.3. Section 6 discusses the boundary contribution using methods for pseudodifferential boundary value problems, here of negative order $-\frac{n+1}{2}$. While such problems have a long history [6], the magnitude problem connects to recent developments for boundary problems for the fractional Laplacians [14, 16]. The resulting boundary contribution to the expansion coefficients $c_{j}$ involves the coefficients $C^{j}$ and the geometry of $\partial X$ inside $X$. Sections 5 and 6 rely on and motivate the purely analytic results for $\mathcal{Z}_{X}$ in [12], which we here exploit for geometric applications.
@@ 3.1. Relation to Euler characteristic
A corresponding connection between the constant term in $\mathcal{M}_{X}$ and the Euler characteristic fails for domains $X\subset\mathbb{R}^{3}$. With $H$ the mean curvature of $\partial X$, $c_{3}(X)=\delta\int_{\partial X}H^{2}$ is a multiple of the Willmore energy by [11]. The Willmore energy, however, can be arbitrarily large on surfaces of genus 00 and is not determined by the Euler characteristic.
@@ 3.5. Relation to residues of manifolds
    $\Gamma(s)f(s)\sim\sum_{j=0}^{\infty}\frac{a_{j}}{\beta_{j}+s}+\sum_{j=0}^{\infty}\frac{A_{j}}{\gamma_{j}-s},$
@@
The assumptions on $e(t)=\langle\mathcal{Z}_{X}(t)1,1\rangle$ from Proposition 3.6 are trivially satisfied; indeed $e(t)=O(t^{-\infty})$ as $t\to\infty$ and $e(t)$ is real analytic near 00. There is therefore, by Proposition 3.6, a close relation between the expectation value of the magnitude operator $\langle\mathcal{Z}_{X}(t)1,1\rangle$ and the residue for an arbitrary metric space $(X,\mathrm{d})$.
@@ 3.6. Computations for cylinders
Assume that $M\subseteq\mathbb{R}^{N}$ is an $n-1$-dimensional compact submanifold. For $T>0$, consider the compact submanifold with boundary
@@
$\mathcal{M}_{X_{T}}(R)\sim$ $\frac{\mathrm{vol}_{n-1}(M)}{n!\omega_{n}}TR^{n}+\frac{(n+1)\mathrm{vol}_{n-1}(M)}{2\cdot n!\omega_{n}}R^{n-1}+$
@@
3.7. Asymptotically polynomial behavior in dimension $3$
@@ 4. Weight distributions and magnitude
In the work [12] the notions of distance functions having property (MR) and (SMR) were introduced. We refer the reader to [12, Section 3] for full details. The reader should keep in mind that (SMR)$\Rightarrow$ (MR) and that both (MR) and (SMR) are inherited by smooth subdomains. Domains and subspace distances in Euclidean space satisfy property (SMR). More generally, a compact subdomain or a subspace $X$ of a manifold $M$ satisfies (SMR) if (SMR) holds on $M$, e.g. if the restriction of $\mathrm{d}^{2}$ to $X\times X$ is smooth. For the geodisic distance this occurs when the diameter of $X$ is smaller than the injectivity radius of $M$. The sphere $S^{n}$ with its geodesic distance has property (MR) but not (SMR) by [12, Proposition 3.6]. Tori and real projective space with their geodesic distances fail to satisfy (MR) in dimension $>1$ by [12, Proposition 3.17]. In the special cases we know that (MR) fails, the magnitude asymptotics can nevertheless be computed by the same formalism as in Theorem 2.1, because Theorem 5.10 below applies.
@@
For the full proof of Theorem 4.6 we refer to [12, Section 4], but we sketch the idea here. The idea in the proof is to use that property $(SMR)$ allows us to replace $\mathcal{Z}_{X}$ with a localization $Q_{X}$ to an operator whose integral kernel is supported near the diagonal. The operator $Q_{X}$ is an elliptic pseudodifferential operator with parameter $R$ of order $-n-1$, we discuss $Q_{X}$ in further detail below in Theorem 5.3. A computation of the principal symbol of $Q_{X}$ shows that $Q_{X}$ is a lower order perturbation of a fractional Laplacian with parameter $(R^{2}+\Delta)^{-(n+1)/2}$ – the Laplacian comes from a Riemannian metric defined from the Taylor expansion of the distance function and the fractional power is defined relative to an extension to an ambient manifold. Therefore item b) follows from the Gårding inequality. Item a) follows from item b) using the meromorphic Fredholm theorem. Item c) is a direct consequence of item b).
@@
Let $X$ be an $n$-dimensional compact manifold with $C^{0}$-boundary and a distance function $\mathrm{d}$ satisfying property (MR) on a sector $\Gamma$ with non-trivial intersection with $[0,\infty)$. Then the magnitude function is given by
@@ 5.1. Localizing to the diagonal
In local coordinates, the full symbol of $Q$ admits a classical asymptotic expansion $\sum_{j=0}^{\infty}q_{j}$ where the $-n-1-j$-homogeneous functions $q_{j}\in C^{\infty}((T^{*}M\oplus\mathbb{C}_{+})\setminus(M\times\{0\}))$ are given by $q_{0}=\sigma_{-n-1}(Q)(x,\xi,R)$ and for $j>0$, $q_{j}$ is in coordinates given by
$q_{j}(x,\xi,R)=$$\begin{cases}&\sum_{\gamma\in I_{j},\mathrm{rk}(\gamma)<(n+1)/2}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)}+\qquad\mbox{for $n$ odd}\\ &-\sum_{\gamma\in I_{j},\mathrm{rk}(\gamma)\geq(n+1)/2}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})\left[(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)}\log(R^{2}+g_{G}(\xi,\xi))\right],\\ {}\\ &\sum_{\gamma\in I_{j}}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)},\ \mbox{for $n$ even.}\end{cases}$
@@
The statement of the theorem is local in nature, so it suffices to compute with $K_{0}$ in local coordinates. Standard considerations show that $K_{0}$ is a conormal distribution $K_{0}\in CI^{-n-1}(Z;\mathrm{Diag}_{M}\times\{0\})$ where $Z=U\times\mathbb{R}$ is a neighborhood of $\mathrm{Diag}_{M}\times\mathbb{R}\subseteq M\times M\times\mathbb{R}$. The arguments in [12, Theorem 2.9] ensures that $Q$ indeed is a pseudodifferential operator with parameter of order $-n-1$ with the prescribed principal symbol.
The computationally delicate issue is that of finding the full symbol in local coordinates. We do so by expanding $K_{0}$ near $x=y$ and inverse transform in $(x-y,\eta)$ to the symbol depending on $(\xi,R)$. Using the Taylor expansion (11), with $v=x-y$, we write
@@
For small $v=x-y$, we can Taylor expand
@@
Let $M$ be an $n$-dimensional compact manifold with a distance function $\mathrm{d}$ regular at the diagonal and smooth off-diagonally, e.g. the subspace distance on a submanifold of Euclidean space. Set $\mu:=(n+1)/2$. Then $\mathcal{Z}$ is an elliptic pseudodifferential operator with parameter $R$ and order $-n-1$, and $\mathcal{Z}-Q$ is smoothing with parameter $R>0$. In particular, for a suitable $R_{0}>0$ and $R>R_{0}$ the operator
@@
is smoothing and exponentially decaying as $R\to+\infty$. We conclude that $Z=Q+L$ is an elliptic pseudodifferential operator with parameter $R$ and order $-n-1$ from Theorem 5.3. Invertibility of $\mathcal{Z}$ for large enough $R$ follows from the ellipticity with parameter of $Q$. And indeed, since $\mathcal{Z}-Q$ is smoothing with parameter, so is $\mathcal{Z}^{-1}-Q^{-1}$ and the norm estimate follows. Therefore,
@@ 5.2. Asymptotic expansions for compact manifolds
    Each $a_{j,0}=a_{j,0}(x,R)$ is homogeneous of degree $m-j$ in $R$.
@@
$n!\omega_{n}\mathcal{M}_{M}(R)=$ $\mathrm{vol}_{n}(M,g)R^{n}+\frac{n+1}{6}\int_{M}s\mathrm{d}xR^{n-1}+$
@@
By well known invariant theory, $c_{j}(M)$ is an integral over $M$ of a universal polynomial in the covariant derivatives of the curvature of $M$ where the total degree is $\leq j$. The only possible invariants in degree 00 is the volume and in degree $2$ it is the integral of the scalar curvature. The universality implies that $c_{0}(M)=\alpha_{0}\mathrm{vol}_{n}(M,g)$ and $c_{2}(M)=\alpha_{2}\int_{M}s\mathrm{d}V$ for dimensional constants $\alpha_{0}$ and $\alpha_{2}$. The values $\alpha_{0}=1$ and $\alpha_{2}=\frac{n+1}{6}$ can be read of from works of Willerton [31, Theorem 11]. Since the Taylor expansion of the geodesic distance function near the boundary only contains even degree terms, $c_{j}(M)=0$ if $j$ is odd. ∎
@@ 6.1. Wiener-Hopf factorization and inverting $\mathcal{Z}_{X}$
    where $Q^{\partial}_{\pm}$ are operators of order $-\mu$. For this factorization to make sense, we need that $Q^{\partial}_{+}$ comes from an operator on $\partial X\times\mathbb{R}$ that preserves support in $\partial X\times[0,\infty)$ and similarly that $Q^{\partial}_{-}$ comes from an operator on $\partial X\times\mathbb{R}$ that preserve supports in $\partial X\times(-\infty,0]$.
@@
Here $h_{0}$ is a smooth function and $h_{\pm}$ are smooth functions (for $\xi^{\prime}\neq 0$) of degree $+1$ in $\xi^{\prime}$. The functions $h_{0}$ and $h_{\pm}$ come from the polynomial factorization
@@
    The symbols $q_{\pm,j}$ extend holomorphically to the half-plane ±$\pm\mathrm{Im}(\xi_{n})>0$ and for $j>0$ can be written as .
@@
    where $b_{\pm,j,k}$ is homogeneous of degree $-k$ in $(\xi^{\prime},R)$.
@@
where $\mathfrak{b}_{\pm,j,k}$ is homogeneous of degree $-k$ in $(\xi^{\prime},R)$ and can be explicitly computed. We now define
@@
where $\mathfrak{w}_{\pm,j,k}$ is homogeneous of degree $-k$ in $(\xi^{\prime},R)$.

### v3

@@ 1.1. Acknowledgements
We thank Magnus Fries, Daniel Grieser, Gerd Grubb and Mark Meckes for useful comments on this article and Tony Carbery, Daniel Grieser, Gerd Grubb, Tom Leinster, Rafe Mazzeo, Mark Meckes, Niels Martin Møller, Grigori Rozenblum, Jan-Philip Solovej and Simon Willerton for fruitful discussions.
@@ 2. Overview and main results
    Analogous to Kac’ famous question [19], one can “magnitude the shape of a drum” for balls: if $X$ is a smooth, compact domain in $\mathbb{R}^{n}$, $B$ is a ball and $\mathcal{M}_{X}(R)=\mathcal{M}_{B}(R)+o(R^{n-1})$, then $X$ is isometric to $B$. See Subsection $3.$
@@
Suppose that $M\subseteq\mathbb{R}^{N}$ is a compact n1n-1-dimensional submanifold, $T>0$ a parameter and $X_{T}=M\times[0,T]\subset\mathbb{R}^{N+1}$ is the cylinder on $M$ of height $T$. The space $X_{T}$ is a compact, $n$-dimensional manifold with boundary, depicted in Figure 1(a). Equipping $X_{T}$ with the subspace distance, Proposition 3.9 relates the magnitude of $X_{T}$ with the magnitude of the base $M$:
@@
Indeed, the volume of $X_{r}$ is $\pi(\sinh(2r)-2r)$, the area of $\partial X_{r}$ is $2\pi(\cosh(2r)-1)$, s=6s=-6 in real hyperbolic $3$-space and $H=H_{0}(r)$ is the mean curvature of the boundary of the ball, which only depends on $r$ by rotational invariance.
@@
where $H_{\mathrm{d}^{2}}(x,\cdot)$ is a Riemannian metric on $X$. In local coordinates $C^{j}$ is a symmetric $j$-form in xyx-y. The Taylor expansion (3) translates into an expansion of the symbol of the pseudodifferential operator $Q_{X}$, see Theorem 5.3, where terms for larger $j$ contribute to terms of the symbol of lower order jointly in $\xi$ and $R$.
@@
The analysis in Section 5 for a closed manifold illustrates the general approach taken in this article. For a manifold with boundary the inverse $Q_{X}(R)^{-1}$ decomposes into the previously studied interior part and a new boundary contribution, as described in Theorem 6.3. Section 6 discusses the boundary contribution using methods for pseudodifferential boundary value problems, here of negative order $\frac{n+1}{2}$. While such problems have a long history [6], the magnitude problem connects to recent developments for boundary problems for the fractional Laplacians [14, 16]. The resulting boundary contribution to the expansion coefficients $c_{j}$ involves the coefficients $C^{j}$ and the geometry of $\partial X$ inside $X$. Sections 5 and 6 rely on and motivate the purely analytic results for $\mathcal{Z}_{X}$ in [12], which we here exploit for geometric applications.
@@ 3.1. Relation to Euler characteristic
A corresponding connection between the constant term in $\mathcal{M}_{X}$ and the Euler characteristic fails for domains $X\subset\mathbb{R}^{3}$. With $H$ the mean curvature of $\partial X$, $c_{3}(X)=\delta\int_{\partial X}H^{2}$ is a multiple of the Willmore energy by [11]. The Willmore energy, however, can be arbitrarily large on surfaces of genus 0 and is not determined by the Euler characteristic.
@@ 3.5. Relation to residues of manifolds
    $\Gamma(s)f(s)\sim\sum_{j=0}^{\infty}\frac{a_{j}}{\beta_{j}+s}+\sum_{j=0}^{\infty}\frac{A_{j}}{\gamma_{j}-s},\$
@@
The assumptions on $e(t)=\langle\mathcal{Z}_{X}(t)1,1\rangle$ from Proposition 3.6 are trivially satisfied; indeed $e(t)=O(t^{-\infty})$ as $t\to\infty$ and $e(t)$ is real analytic near 0. There is therefore, by Proposition 3.6, a close relation between the expectation value of the magnitude operator $\langle\mathcal{Z}_{X}(t)1,1\rangle$ and the residue for an arbitrary metric space $(X,\mathrm{d})$.
@@ 3.6. Computations for cylinders
Assume that $M\subseteq\mathbb{R}^{N}$ is an n1n-1-dimensional compact submanifold. For $T>0$, consider the compact submanifold with boundary
@@
$\mathcal{M}_{X_{T}}(R)\sim\$$\frac{\mathrm{vol}_{n-1}(M)}{n!\omega_{n}}TR^{n}+\frac{(n+1)\mathrm{vol}_{n-1}(M)}{2\cdot n!\omega_{n}}R^{n-1}+$
@@
3.7. Asymptotically polynomial behavior in dimension 33
@@ 4. Weight distributions and magnitude
[…]
@@
For the full proof of Theorem 4.6 we refer to [12, Section 4], but we sketch the idea here. The idea in the proof is to use that property $(SMR)$ allows us to replace $\mathcal{Z}_{X}$ with a localization $Q_{X}$ to an operator whose integral kernel is supported near the diagonal. The operator $Q_{X}$ is an elliptic pseudodifferential operator with parameter $R$ of order n1-n-1, we discuss $Q_{X}$ in further detail below in Theorem 5.3. A computation of the principal symbol of $Q_{X}$ shows that $Q_{X}$ is a lower order perturbation of a fractional Laplacian with parameter $(R^{2}+\Delta)^{-(n+1)/2}$ – the Laplacian comes from a Riemannian metric defined from the Taylor expansion of the distance function and the fractional power is defined relative to an extension to an ambient manifold. Therefore item b) follows from the Gårding inequality. Item a) follows from item b) using the meromorphic Fredholm theorem. Item c) is a direct consequence of item b).
@@
Let $X$ be an $n$-dimensional compact manifold with $C^{0}$-boundary and a distance function $\mathrm{d}$ satisfying property (MR) on a sector $\Gamma$ with non-trivial intersection with $0,\infty)$. Then the magnitude function is given by
@@ 5.1. Localizing to the diagonal
In local coordinates, the full symbol of $Q$ admits a classical asymptotic expansion $\sum_{j=0}^{\infty}q_{j}$ where the n1j-n-1-j-homogeneous functions $q_{j}\in C^{\infty}((T^{*}M\oplus\mathbb{C}_{+})\setminus(M\times\{0\}))$ are given by $q_{0}=\sigma_{-n-1}(Q)(x,\xi,R)$ and for $j>0$, $q_{j}$ is in coordinates given by
$q_{j}(x,\xi,R)=$ $\begin{cases}&\sum_{\gamma\in I_{j},\mathrm{rk}(\gamma)<(n+1)/2}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)}+\hskip 18.49988pt\mbox{for $n$ odd}\\ &-\sum_{\gamma\in I_{j},\mathrm{rk}(\gamma)\geq(n+1)/2}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})\left[(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)}\log(R^{2}+g_{G}(\xi,\xi))\right],\\ {}\\ &\sum_{\gamma\in I_{j}}\mathfrak{c}_{\mathrm{rk}(\gamma),n}C^{(\gamma)}(x,-D_{\xi})(R^{2}+g_{\mathrm{d}}(\xi,\xi))^{-(n+1)/2+\mathrm{rk}(\gamma)},\ \mbox{for $n$ even.}\end{cases}$
@@
The statement of the theorem is local in nature, so it suffices to compute with $K_{0}$ in local coordinates. Standard considerations show that $K_{0}$ is a conormal distribution $K_{0}\in CI^{-n-1}(Z;\mathrm{Diag}_{M}\times\{0\})$ where $Z=U\times\mathbb{R}$ is a neighborhood of $\mathrm{Diag}_{M}\times\mathbb{R}\subseteq M\times M\times\mathbb{R}$. The arguments in [12, Theorem 2.9] ensures that $Q$ indeed is a pseudodifferential operator with parameter of order n1-n-1 with the prescribed principal symbol.
The computationally delicate issue is that of finding the full symbol in local coordinates. We do so by expanding $K_{0}$ near $x=y$ and inverse transform in $(x-y,\eta)$ to the symbol depending on $(\xi,R)$. Using the Taylor expansion (11), with v=xyv=x-y, we write
@@
For small v=xyv=x-y, we can Taylor expand
@@
Let $M$ be an $n$-dimensional compact manifold with a distance function $\mathrm{d}$ regular at the diagonal and smooth off-diagonally, e.g. the subspace distance on a submanifold of Euclidean space. Set $\mu:=(n+1)/2$. Then $\mathcal{Z}$ is an elliptic pseudodifferential operator with parameter $R$ and order n1-n-1, and $\mathcal{Z}-Q$ is smoothing with parameter $R>0$. In particular, for a suitable $R_{0}>0$ and $R>R_{0}$ the operator
@@
is smoothing and exponentially decaying as $R\to+\infty$. We conclude that $Z=Q+L$ is an elliptic pseudodifferential operator with parameter $R$ and order n1-n-1 from Theorem 5.3. Invertibility of $\mathcal{Z}$ for large enough $R$ follows from the ellipticity with parameter of $Q$. And indeed, since $\mathcal{Z}-Q$ is smoothing with parameter, so is $\mathcal{Z}^{-1}-Q^{-1}$ and the norm estimate follows. Therefore,
@@ 5.2. Asymptotic expansions for compact manifolds
    Each $a_{j,0}=a_{j,0}(x,R)$ is homogeneous of degree mjm-j in $R$.
@@
$n!\omega_{n}\mathcal{M}_{M}(R)=$ $\mathrm{vol}_{n}(M,g)R^{n}+\frac{n+1}{6}\int_{M}s\mathrm{d}xR^{n-2}+$
@@
By well known invariant theory, $c_{j}(M)$ is an integral over $M$ of a universal polynomial in the covariant derivatives of the curvature of $M$ where the total degree is $\leq j$. The only possible invariants in degree 0 is the volume and in degree $2$ it is the integral of the scalar curvature. The universality implies that $c_{0}(M)=\alpha_{0}\mathrm{vol}_{n}(M,g)$ and $c_{2}(M)=\alpha_{2}\int_{M}s\mathrm{d}V$ for dimensional constants $\alpha_{0}$ and $\alpha_{2}$. The values $\alpha_{0}=1$ and $\alpha_{2}=\frac{n+1}{6}$ can be read of from works of Willerton [31, Theorem 11]. Since the Taylor expansion of the geodesic distance function near the boundary only contains even degree terms, $c_{j}(M)=0$ if $j$ is odd. ∎
@@ 6.1. Wiener-Hopf factorization and inverting $\mathcal{Z}_{X}$
    where $Q^{\partial}_{\pm}$ are operators of order $\mu$. For this factorization to make sense, we need that $Q^{\partial}_{+}$ comes from an operator on $\partial X\times\mathbb{R}$ that preserves support in $\partial X\times[0,\infty)$ and similarly that $Q^{\partial}_{-}$ comes from an operator on $\partial X\times\mathbb{R}$ that preserve supports in $\partial X\times(-\infty,0]$.
@@
Here $h_{0}$ is a smooth function and $h_{\pm}$ are smooth functions (for $\xi^{\prime}\neq 0$) of degree +1+1 in $\xi^{\prime}$. The functions $h_{0}$ and $h_{\pm}$ come from the polynomial factorization
@@
    The symbols $q_{\pm,j}$ extend holomorphically to the half-plane $\pm\mathrm{Im}(\xi_{n})>0$ and for $j>0$ can be written as .
@@
    where $b_{\pm,j,k}$ is homogeneous of degree k-k in $(\xi^{\prime},R)$.
@@
where $\mathfrak{b}_{\pm,j,k}$ is homogeneous of degree k-k in $(\xi^{\prime},R)$ and can be explicitly computed. We now define
@@
where $\mathfrak{w}_{\pm,j,k}$ is homogeneous of degree k-k in $(\xi^{\prime},R)$.